Two cars leave town at the same time going in opposite directions. one travels at 35 km/h and the other at 55 km/h. in how many hours will they be 200 km apart?
Answer by pushpaharan(47)

let x be the time taken for 200 km apeart.
in x hrs car "A" will go x*35 km (distance =speed * time)
in x hrs car "B" will go x*55 km
togather both will go 35x+55x = 200
90x = 200
x=200/90 hts
x= 2 hrs and 2/9 hrs
= 2 hrs and 2/9*60 mins
= 2 hrs and 13 mins